Grade upto college level Physical Chemistry

In the mineral, spinel, having the formula MgAl2O4, oxide irons are arranged in the cubic close packing. Mg2+ions occupy the tetrahedral voids while Al3+ions occupy the octahedral voids. (i) What is the percentage of tetrahedral voids occupied by Mg2+ions? (ii) What is the percentage of octahedral voids occupied by Al3+ions?

aditya kashyap
12 Years agoGrade upto college level

3 Answers

Kshitiz
9 Years ago
12.5% tetrahedral voids and 50% octahedral voids gehsjjsksksjdhdkkshshsjjshshshshshhsjsjshshdhjskshshsujshdhdududjjsjsjdjdjhdhsjsjsjsisidjjdjdhdjdjdjdhhdhdjdjdjdjhfhfhfjjdndhdhdjjdjdjdjd
Nitin
8 Years ago
According to formula MgAl2O4 oxygen contains 4 atoms that means it is forming CCP or FCC also there is only 1 mg atom and 2 Al atom.Z=4Octahedral voids=4Tetrahedral voids=8As mg is occupying tetrahedral voids so percentage = (1/8)×100=12.5%.And percentage of octahedral voids occupied by Al atoms=(2/4)×100=50%.
Amitesh Kumar Singh
4 Years ago
In the mineral spinel, MgAl 
2
 O 
4
   oxide ions formed ccp lattice Mg 
+2
  ions occupy tetrahedral voids while Al 
+3
  ions occupy the octahedral voids then the percentage of tetrahedral, octahedral voids occupied by Mg 
+2
  and Al 
+3
  ions respectively
Then find 0.1(3p+4q)